40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 24 23:50:35 2012, Koji, Update, IOO, MC WFS refusing to work 
    Reply  Thu Nov 29 00:15:13 2012, Den, Koji, Update, IOO, MC WFS work wfs_fb.pdfDSC_4942.JPG
Message ID: 7761     Entry time: Thu Nov 29 00:15:13 2012     In reply to: 7741
Author: Den, Koji 
Type: Update 
Category: IOO 
Subject: MC WFS work 

Quote:

The instability started about 48hour ago, that means my work on the AP table did not 
made immediate trouble. But it does not mean anything.

For now, the WFS outputs are off. More work is needed to find what's wrong.

 

 The problem was caused by low reflectivity of the mirror that splits MC reflected beam into two: first goes to trash, second - to WFS. Power before the mirror was 100mW, reflected beam that goes to WFS was 0.3mW. Using dataviewer we learnt that the beam intensity was ~5 times more in the past.

This happened because the mirror position was adjusted a few days ago. Its reflection depends on the angle of incidence and amount of light to WFS was significantly reduced. We could either increase the angle of incidence or use two mirrors with high reflectivity instead of this with high transmission.

We've chosen the second variant not to confuse anyone in future with non-45 degrees angles. We are now using one mirror with reflectivity 98% to direct most power to the trash while other 2% are directed using the second mirror to WFS path. We now have 0.7 mW on WFS1 and 1.3 mW on WFS2.

Then we adjusted WFS 

  • blocked the beam and run scripts/MC/WFS/WFS_FilterBank_offsets to calculate offsets in the WFS servo
  • aligned MC and centered beams on WFS 1 and 2
  • provided excitation to MC1 at 5 Hz (400 counts) and adjusted I&Q phase rotation
  • adjusted the gain and changed it in MC autolocker (reduced from 0.25 to 0.15 as we now have more power of WFS as before)

We were able to close the loops. The phase margin is too low though, we need to improve feedback filters.

DSC_4942.JPG

Attachment 1: wfs_fb.pdf  50 kB  | Hide | Hide all
wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf
ELOG V3.1.3-